OpenStudy (anonymous):

Would the simplified version of h(x)=(x^2-3x-4)/(x-4) be x+1? If so, how would I simplify h(x) and write the new function as g(x)?

OpenStudy (mertsj):

\[g(x)=\frac{h(x)}{x-4}\]

OpenStudy (anonymous):

even though it says to simplify h(x) first and then write the new function as g(x)? @Mertsj

OpenStudy (mertsj):

Guess you should have posted the correct problem.

OpenStudy (anonymous):

I thought I did? My original question was "Would the simplified version of h(x)=(x^2-3x-4)/(x-4) be x+1? If so, how would I simplify h(x) and write the new function as g(x)?" .

OpenStudy (mertsj):

Yes. The simplified version is x+1 Write g(x)=x+1

OpenStudy (anonymous):

sorry if there was miscommunication

OpenStudy (anonymous):

thank you @Mertsj, do you know how I should approach the question "What signficant graphical feature is produced when a function like h(x) can be simplified to g(x)?"

OpenStudy (mertsj):

I think a hole in the graph.

OpenStudy (anonymous):

okay, thank you for all of your help! :)

OpenStudy (mertsj):

yw